OpenStudy (anonymous):

How do I do this? A line has a slope of –2 and passes through the point (4, –9). Which is the equation of the line? y = –2x – 17 y = –2x – 1 y = –2x + 1 y = –2x + 17

OpenStudy (anonymous):

@iGreen @IMStuck

OpenStudy (anonymous):

by equation y-y1=m(x-x1)

OpenStudy (anonymous):

where x1 and y1 are 4 and -9 respectively

OpenStudy (anonymous):

and m is slope that is -2

OpenStudy (anonymous):

Y=mx+b m=slope Y=-2x+b (x,y) (4,-9) sub in 4 where x is and -9 where y is (-9)=-2(4)+b -9=-8+b Add 8 on both sides y=-2x-1

OpenStudy (anonymous):

so b option is correct

OpenStudy (anonymous):

Oh, I see. :) It's a lot easier when you can see it like that. Thanks!!

OpenStudy (anonymous):

yeah no problem!